1/1

Heinrich Heine,

The Lorelei

tradukita de L. W. Garnham

I do not know what it signifies. That I am so sorrowful? A fable of old Times so terrifies, Leaves my heart so thoughtful.

The air is cool and it darkens, And calmly flows the Rhine; The summit of the mountain hearkens In evening sunshine line.

The most beautiful Maiden entrances Above wonderfully there, Her beautiful golden attire glances, She combs her golden hair.

With golden comb so lustrous, And thereby a song sings, It has a tone so wondrous, That powerful melody rings.

The shipper in the little ship It effects with woe sad might; He does not see the rocky slip, He only regards dreaded height.

I believe the turbulent waves Swallow the last shipper and boat; She with her singing craves All to visit her magic moat.

Traduko de la Germana poemo “Die Lorelei” de Heinrich Heine (∗1797-12-13 – †1856-02-17) en la Anglan de L. W. Garnham.

Arg-2-14 (2003-10-11 23:04:57)

L.W. Garnham, Bachelor of Arts, LEGENDS OF THE RHINE; mentioned in: A Tramp Abroad. Vol 1-2. Leibzig: Tauchnitz, 1880 Band I, Mark Twain 1880

http://www.poezio.net/version?poem-id=2&version-id=,14